# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

### v1

@@ 3 Error Analysis of the SGIP method for RDA Equation
$\varepsilon_{\mathrm{react}}\leq C_{5}\Delta t^{p}.$ (55)
Lemma 5 (Resampling error).
Under Assumption 1, the resampling error satisfies:
$\mathbb{E}[\varepsilon_{\mathrm{resamp}}]=\mathbb{E}[\|\widetilde{u}_{n+1}-\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}]\leq C_{6}(N\Delta x^{d})^{-1/2}$ (56)
where the constant $C_{6}$ depends on the domain size $L$, dimension $d$, and the total mass $M$.
Proof.
The resampling step converts the continuous density field $\widetilde{u}^{\mathrm{R}}_{n+1}$ back to a particle representation. This process introduces statistical sampling error:
The resampling step generates $N$ new particles according to the multinomial distribution with probabilities:
$p_{j}=\frac{\widetilde{u}^{\mathrm{R}}_{n+1,j}\cdot(\Delta x)^{d}}{M}$ (57)
where $M=\sum_{j=1}^{M^{d}}\widetilde{u}^{\mathrm{R}}_{n+1,j}\cdot(\Delta x)^{d}$ is the total mass.
Let $n_{j}$ be the number of particles assigned to bin $j$ with $\sum_{j=1}^{M^{d}}n_{j}=N$. The new density estimate after resampling is:
$\widetilde{u}_{n+1,j}=\frac{M\cdot n_{j}}{N(\Delta x)^{d}}$ (58)
The expected value is:
$\mathbb{E}[\widetilde{u}_{n+1,j}]=\frac{M}{N(\Delta x)^{d}}\cdot\mathbb{E}[n_{j}]=\frac{M}{N(\Delta x)^{d}}\cdot Np_{j}=\widetilde{u}^{\mathrm{R}}_{n+1,j}$ (59)
so the resampling is unbiased.
The variance is:
$\mathrm{Var}(\widetilde{u}_{n+1,j})=\left(\frac{M}{N(\Delta x)^{d}}\right)^{2}\cdot\mathrm{Var}(n_{j})=\left(\frac{M}{N(\Delta x)^{d}}\right)^{2}\cdot Np_{j}(1-p_{j})$ (60)
Using $p_{j}\leq\frac{\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(\Delta x)^{d}}{M}$, we get:
$\mathrm{Var}(\widetilde{u}_{n+1,j})\leq\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}}{N(\Delta x)^{d}}$ (61)
Summing over all bins:
$\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\widetilde{u}_{n+1,j}-\widetilde{u}^{\mathrm{R}}_{n+1,j}|^{2}\right]$ $\leq\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}}{N(\Delta x)^{d}}$
$\leq\frac{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(2L)^{d}}{N(\Delta x)^{d}}$
Taking square roots:
$\mathbb{E}[\|\widetilde{u}_{n+1}-\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{2}}]\leq\frac{C}{\sqrt{N}\Delta x^{d/2}}=C_{6}(N\Delta x^{d})^{-1/2}$ (62)
where $C_{6}=\sqrt{M\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}(2L)^{d}}$.
Under Assumption 1, $M$ and $\|\widetilde{u}^{\mathrm{R}}_{n+1}\|_{L^{\infty}}$ are uniformly bounded, ensuring $C_{6}$ is finite. ∎
Now, we are ready to prove Theorem 3.
Proof of Theorem 3. From Eqs.(15)(3), and all the lemmas above, we can obtain:
@@
$\mathbb{E}\,\|\widetilde{u}_{n+1}-u_{n+1}\|_{L^{2}}$ $\leq\mathbb{E}[\varepsilon_{\mathrm{propag}}]+\mathbb{E}[\varepsilon_{\mathrm{dens}}]+\mathbb{E}[\varepsilon_{\mathrm{react}}]+\mathbb{E}[\varepsilon_{\mathrm{resamp}}]+\mathbb{E}[\varepsilon_{\mathrm{split}}]$
$\leq(1+C_{1}\Delta t)\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}+C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$
$\quad+C_{5}\Delta t^{p}+C_{6}(N\Delta x^{d})^{-1/2}+C_{2}\Delta t^{2}$ (63)
For the optimal case where we use exact reaction integration ($p\to\infty$) or second-order methods ($p=2$), we obtain:
$\mathbb{E}\,\|\widetilde{u}_{n+1}-u_{n+1}\|_{L^{2}}\leq(1+C_{1}\Delta t)\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}+C_{4}\Delta x+C_{7}\Delta t^{2}+C_{8}(N\Delta x^{d})^{-1/2}$ (64)
where $C_{7}=C_{2}+C_{4}$, $C_{8}=C_{3}+C_{6}$.
Since the initial condition is sampled exactly from $u_{0}$, we have $\varepsilon_{0}=0$. By the discrete Gronwall’s inequality, for $n\Delta t\leq T$:
@@
$\mathbb{E}\,\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ $\leq\frac{e^{C_{1}T}-1}{C_{1}}\left(C_{7}(N\Delta x^{d}\Delta t)^{-1/2}+C_{8}\frac{\Delta x}{\Delta t}+C_{2}\Delta t\right)$
$\leq C_{T}\left(\Delta t+(N\Delta x^{d}\Delta t)^{-1/2}+\frac{\Delta x}{\Delta t}\right)$ (65)
@@
where $C_{T}=\frac{e^{C_{1}T}-1}{C_{1}}\max(C_{7},C_{8},C_{2})$. This completes the proof of Theorem 3.
@@ 4 Numerical Experiments
All experiments are conducted on the computational domain $\Omega=[-L,L]$ with $L=60$. The initial condition is given by a uniform distribution on the interval $[0,1]$:
$u_{0}(x)=\chi_{[0,1]}(x)=\begin{cases}1&\text{for }x\in[0,1]\\ 0&\text{otherwise}\end{cases}$ (66)
@@
For the reference FDM simulations, we employ a high-resolution discretization with spatial step $\delta x=10^{-3}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions. For the SGIP, we use $N=10^{6}$ particles to approximate the density distribution, with the spatial domain discretized into $M=150$ bins for density estimation, corresponding to a bin width of $\Delta x=0.8$. The time step is set to $\Delta t=0.5$.
First, we consider the classical FKPP equation:
$u_{t}=u_{xx}+u(1-u).$ (67)
(a) t=5
(b) t=10
(c) t=15
(d) t=20
@@
Figure 1 displays the numerical solutions of the standard FKPP equation at different time instances. It can be observed that the SGIP method results show excellent agreement with FDM. As expected for the FKPP equation, the shape of the wave front remains unchanged during propagation, demonstrating the characteristic traveling wave structure.
Next, we consider a reaction-diffusion equation with cubic nonlinear reaction term:
$u_{t}=u_{xx}+u^{2}(1-u).$ (68)
(a) t=5
(b) t=10
(c) t=15
(d) t=20
@@
Figure 2 shows the numerical results for the case with cubic nonlinear reaction term. Notably, the cubic nonlinearity $u^{2}(1-u)$ leads to a visibly slower propagation speed compared to the standard FKPP case, while maintaining a stable wave profile. The SGIP method still accurately captures the wave front propagation behavior and maintains consistency with FDM results, demonstrating the robustness of the SGIP method when dealing with more complex nonlinear reaction terms.
Finally, we consider a reaction-diffusion equation with Arrhenius-type reaction term:
$u_{t}=u_{xx}+e^{-E/u}(1-u)$ (69)
where we choose $E=\frac{1}{2}$.
(a) t=5
(b) t=10
(c) t=15
(d) t=20
@@
Figure 3 presents the numerical results for the exponential reaction term case. This reaction term possesses singularity near $u=0$, posing additional challenges for numerical methods. The SGIP method demonstrates satisfactory performance in capturing the wave propagation under this singular nonlinearity. It is observed that the wave front propagates at a notably slower speed compared to the standard FKPP case, while maintaining a coherent wave profile. These results further validate the robustness of the SGIP in handling equations with complex nonlinearities.
To quantitatively compare the effects of different reaction terms on propagation speed, we plot the evolution of wave front position over time for all three cases in Fi gure 4. These numerical experiments comprehensively demonstrate the effectiveness and accuracy of the SGIP method for solving various types of reaction-diffusion equations.
@@
We now extend our investigation to 2D RAD systems to further validate the SGIP method’s capability in handling more complex spatial dynamics. The 2D experiments provide insights into pattern formation, wave propagation in multiple dimensions, and the interaction between diffusion and advection processes.
All two-dimensional simulations are conducted on the computational domain $\Omega=[-60,60]^{2}$. The initial condition is specified by sampling particle positions uniformly from the unit square: $\mathbf{X}_{0}^{i}\sim\text{Uniform}([0,1]^{2})$ for $i=1,\dots,N$, corresponding to an initial density $u_{0}(x,y)=\chi_{[0,1]^{2}}(x,y)$.
For the reference FDM simulations, we employ a high-resolution uniform grid with spatial discretization $\delta x=\delta y=10^{-2}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions.
For the SGIP method, we use $N=3\times 10^{6}$ particles to approximate the density distribution. The spatial domain is discretized into $150^{2}$ uniform bins for density estimation, and the time step is set to $\Delta t=0.5$.
@@
Figure 6: Propagation speed comparison for particle and FDM
Figure 5 demonstrates the evolution of the 2D FKPP equation with uniform initial conditions, showing contour lines at $u=0.1$, $0.5$, and $0.9$. The SGIP method successfully preserves the circular symmetry of the expanding wave front, maintaining rotational invariance in the contour morphology across all three concentration levels. Quantitative comparison in Figure 6 assesses the radial propagation speeds by tracking the intersection points of the $u=0.2$ contour with the x-axis for both SGIP and FDM. The excellent agreement between methods, evidenced by the linear position-time relationship, confirms the constant speed propagation characteristic of FKPP-type equations and validates the accuracy of our SGIP method in capturing two-dimensional front dynamics.
@@
Shear Flow:
$u_{t}=D\Delta u+(\sin y,0)\cdot\nabla u+u(1-u).$ (70)
@@
Figure 8: SGIP vs. FDM on cellular flow
Cat’s Eye Flow:
$u_{t}=D\Delta u+[(-\sin x\cos y,\cos x\sin y)+\delta\cdot(\cos x\sin y,-\sin x\cos y)]\cdot\nabla u+u(1-u),$ (72)
here we set $\delta=2$.
@@
Figure 9: SGIP vs. FDM on cat’s eye flow
Figure 7 demonstrates the effect of a sinusoidal shear flow on the reaction-diffusion dynamics, where the SGIP accurately captures the formation of characteristic shear-aligned patterns, showing how the flow stretches and orients the chemical waves along the shear direction. In Figure 8, the interaction with cellular flow patterns is shown, with the Lagrangian formulation naturally resolving the complex recirculation zones and their impact on front morphology. The cellular flow organizes the reaction fronts into approximately circular structures confined within individual flow cells, while inducing characteristic corrugations along the intercellular boundaries due to hyperbolic stagnation points and associated stretching fields. Figure 9 examines a more intricate cat eye flow combining cellular and shear components, where the SGIP successfully reproduces the competition between diffusion and advection across different diffusion parameter regimes ($D=0.5$ and $D=1$). The method captures the characteristic lobe structures formed by alternating elliptic islands and hyperbolic channels, with lower diffusivity ($D=0.5$) producing sharp lobe boundaries and higher diffusivity ($D=1$) enhancing cross-channel mixing and smoothing front morphology. The SGIP resolves how the flow superposition modulates front propagation through the competing recirculation and transport regions.
4.3 3D Numerical Experiments on ABC Flow
We now extend our investigation to 3D systems with the ABC flow, described by:
$u_{t}=D\Delta u+(A\sin z+C\cos y,B\sin x+A\cos z,C\sin y+B\cos x)\cdot\nabla u+u(1-u)$ (73)
@@
The SGIP employs $5\times 10^{6}$ particles with $100^{3}$ spatial bins and time step $\Delta t=0.5$. In contrast, the FDM requires significantly finer discretization for numerical stability: time step $\delta t=10^{-3}$ and spatial grid of at least $300^{3}$ points, resulting in substantially higher computational cost.
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 10: SGIP vs. FDM on ABC flow with $D=1$
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 11: SGIP vs. FDM on ABC flow with $D=0.5$
@@
Figures 10 and 11 compare both methods on cross-sections through the coordinate planes for diffusion coefficients $D=1$ and $D=0.5$, showing excellent agreement. As $D$ decreases to 0.1 (Figure 12), advection and reaction dominate the dynamics, requiring even finer FDM discretization ($400^{3}$ grid points) that becomes computationally prohibitive due to excessive memory requirements on conventional workstations. The SGIP method, however, remains feasible and accurately captures the complex 3D flow structures across all diffusion regimes.
We further validate the SGIP using more complex nonlinear reaction terms (cubic, Arrhenius), maintaining the same computational parameters ($5\times 10^{6}$ particles, $100^{3}$ bins, $\Delta t=0.5$). However, since these new reaction terms result in markedly different reaction kinetics, the previously used unit ball initial condition would produce dynamics that are too confined and difficult to visualize effectively within the computational domain. To compensate for this and to generate clearly observable front propagation or pattern formation, the initial condition is modified to $u_{0}(\mathbf{x})=\chi_{B(0,5)}(\mathbf{x})$ on a larger support.
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 13: SGIP vs. FDM on ABC Flow with Reaction Term: $u^{2}(1-u)$
(a) SGIP with projection at x=0 plane
(b) FDM with projection at x=0 plane
(c) SGIP with projection at y=0 plane
(d) FDM with projection at y=0 plane
(e) SGIP with projection at z=0 plane
(f) FDM with projection at z=0 plane
Figure 14: SGIP vs. FDM on ABC Flow with Reaction Term: $e^{-\frac{1}{2u}}(1-u)$

### v2

@@ 3 Error Analysis of the SGIP method for RDA Equation
(55) $\varepsilon_{\mathrm{react}}^{2}$ $=\int_{\Omega}|\Psi_{\Delta t}(\tilde{u}^{D}(\mathbf{x}))-S_{\mathcal{R}}(\Delta t)\tilde{u}^{D}(\mathbf{x})|^{2}d\mathbf{x}\leq|\Omega|C_{r}^{2}(\Delta t)^{2(p+1)}.$
Taking the square root, we obtain:
(56) $\varepsilon_{\mathrm{react}}\leq C_{5}(\Delta t)^{p+1}.$
where the constant $C_{5}=\sqrt{|\Omega|}C_{r}$.
Lemma 3.13 (Resampling error).
Under Assumption 1, the resampling step introduces a mean-zero stochastic error with variance satisfying:
(57) $\mathbb{E}[\|\varepsilon_{\mathrm{resamp}}\|^{2}]=\mathbb{E}[\|\widetilde{u}_{n+1}-\widetilde{u}^{\mathrm{R}}_{n+1}\|^{2}]\leq\frac{C_{6}}{N(\Delta x)^{d}},$
where the constant $C_{6}$ depends on the domain size and dimension.
Proof 3.14.
Let $w_{j}=\tilde{u}^{\mathrm{R}}_{n+1,j}$ be the density value in bin $j$ before resampling. The resampling process draws $N$ particles based on the mass distribution. The probability of a new particle falling into bin $j$ is:
(58) $q_{j}=\frac{w_{j}(\Delta x)^{d}}{M},\quad\text{where }M=\sum_{k}w_{k}(\Delta x)^{d}.$
Let $K_{j}$ be the number of particles in bin $j$ after resampling. The vector $(K_{1},\dots,K_{|\mathcal{G}|})$ follows a multinomial distribution with parameters $N$ and probabilities $\{q_{j}\}$. The resampled density is $\tilde{u}_{n+1,j}=\frac{MK_{j}}{N(\Delta x)^{d}}$.
Since $\mathbb{E}[K_{j}]=Nq_{j}$, we have:
(59) $\mathbb{E}[\tilde{u}_{n+1,j}]=\frac{M}{N(\Delta x)^{d}}(Nq_{j})=\frac{M}{(\Delta x)^{d}}\frac{w_{j}(\Delta x)^{d}}{M}=w_{j}.$
Thus, $\mathbb{E}[\tilde{u}_{n+1}]=\tilde{u}^{\mathrm{R}}_{n+1}$.
The variance of the count in bin $j$ is $\mathrm{Var}(K_{j})=Nq_{j}(1-q_{j})$. The expected squared $L^{2}$ error is:
(60) $\mathbb{E}[\|\tilde{u}_{n+1}-\tilde{u}^{\mathrm{R}}_{n+1}\|^{2}]=\frac{M^{2}}{N(\Delta x)^{d}}\sum_{j\in\mathcal{G}}q_{j}(1-q_{j})\leq\frac{C_{6}}{N(\Delta x)^{d}}.$
where the constant $C_{6}=(C_{M})^{2}$, consistent with the mass bound $C_{M}$ in Lemma 3.3.
Now, we are ready to prove Theorem 3.1.
Proof of Theorem 3.1. Recall the decomposition of the global error into the dynamic discrete error and the static projection error:
(61) $\|\widetilde{u}_{n}-u_{n}\|\leq\|\widetilde{u}_{n}-\bar{u}_{n}\|+\|\bar{u}_{n}-u_{n}\|.$
The static projection error is bounded by $\|\bar{u}_{n}-u_{n}\|\leq C_{7}\Delta x$. We define the mean squared dynamic error as $E_{n}=\mathbb{E}[\|\widetilde{u}_{n}-\bar{u}_{n}\|^{2}]$.
Using the error decomposition in Eq.(22) and introducing the ideal discrete operator $\Phi_{h}$, the error evolution at step $(n+1)$ is:
(62) $e_{n+1}=\widetilde{u}_{n+1}-\bar{u}_{n+1}=(\Phi_{h}(\widetilde{u}_{n})-\Phi_{h}(\bar{u}_{n}))+\varepsilon_{\mathrm{consist}}+\varepsilon_{\mathrm{num}}.$
Crucially, from Eq.(23), the stochastic numerical error $\varepsilon_{\mathrm{dens}}+\varepsilon_{\mathrm{resamp}}$ (comprising density estimation variance and resampling noise) has zero mean conditional on the state at time $t_{n}$. Therefore, the cross-term in the squared norm expansion vanishes in expectation:
$E_{n+1}$ $=\mathbb{E}\|\Phi_{h}(\widetilde{u}_{n})-\Phi_{h}(\bar{u}_{n})+\varepsilon_{\mathrm{consist}}+\varepsilon_{\mathrm{num}}\|^{2}$
@@
(63) $=\mathbb{E}\|\Phi_{h}(\widetilde{u}_{n})-\Phi_{h}(\bar{u}_{n})+\varepsilon_{\mathrm{consist}}+\varepsilon_{\mathrm{react}}\|^{2}+\mathbb{E}\|\varepsilon_{\mathrm{dens}}+\varepsilon_{\mathrm{resamp}}\|^{2}.$
We combine the consistency error from Lemma 3.7, the reaction error from Lemma 3.11. At this stage, we assume a practical first-order time integration scheme ($p=1$) is used for the reaction step, which yields $\|\varepsilon_{\mathrm{react}}\|\leq C_{5}(\Delta t)^{2}$. Using the inequality $\|a+b\|^{2}\leq(1+\Delta t)\|a\|^{2}+(1+\frac{1}{\Delta t})\|b\|^{2}$, Lemma 3.5, and Lemma 3.7, we obtain:
$\mathbb{E}\|\Phi_{h}(\widetilde{u}_{n})-\Phi_{h}(\bar{u}_{n})+\varepsilon_{\mathrm{consist}}+\varepsilon_{\mathrm{react}}\|^{2}$
@@
$\leq(1+\Delta t)\mathbb{E}\|\Phi_{h}(\widetilde{u}_{n})-\Phi_{h}(\bar{u}_{n})\|^{2}+\left(1+\frac{1}{\Delta t}\right)\|\varepsilon_{\mathrm{consist}}+\varepsilon_{\mathrm{react}}\|^{2}$
$\leq(1+\Delta t)(1+C_{1}\Delta t)^{2}E_{n}+\frac{2}{\Delta t}\left((C_{2}+C_{5})(\Delta t)^{2}+C_{3}\Delta t\Delta x\right)^{2}$
(64) $\leq(1+C_{8}\Delta t)E_{n}+C_{9}\Delta t(\Delta t+\Delta x)^{2},$
where we have simplified terms for small $\Delta t$.
Combining Lemma 3.9 and Lemma 3.13, the total variance is:
(65) $\mathbb{E}\|\varepsilon_{\mathrm{dens}}+\varepsilon_{\mathrm{resamp}}\|^{2}\leq\frac{C_{4}+C_{6}}{N(\Delta x)^{d}}=\frac{C_{10}}{N(\Delta x)^{d}}.$
Substituting these bounds back into the recurrence for $E_{n}$:
(66) $E_{n+1}\leq(1+C_{8}\Delta t)E_{n}+C_{9}\Delta t(\Delta t+\Delta x)^{2}+\frac{C_{10}}{N(\Delta x)^{d}}.$
Since the initial condition is sampled such that the initial bin masses match the projection, $E_{0}\approx 0$ (or is bounded by initial sampling variance $O(1/N)$). By the discrete Gronwall’s inequality, for $n\Delta t\leq T$:
(67) $E_{n}\leq e^{C_{8}T}\left(C_{9}(\Delta t+\Delta x)^{2}+\frac{C_{10}}{N(\Delta x)^{d}\Delta t}\right).$
Finally, taking the square root to obtain the $L^{2}$-norm estimate, and adding the static projection error $\|\bar{u}_{n}-u_{n}\|\leq C_{10}\Delta x$:
(68) $\mathbb{E}\|\widetilde{u}_{n}-u_{n}\|$ $\leq\sqrt{E_{n}}+C_{7}\Delta x\leq C_{T}\left(\Delta t+\frac{1}{\sqrt{N(\Delta x)^{d}\Delta t}}+\Delta x\right).$
@@
where $C_{T}=e^{\frac{1}{2}C_{8}T}\max\left\{\sqrt{C_{9}},\sqrt{C_{10}}\right\}+C_{7}$. This completes the proof of Theorem 3.1.
@@ 4 Numerical Experiments
This section presents a series of numerical experiments designed to validate the accuracy, robustness, and computational efficiency of the proposed SGIP method. The validation begins with nonlinear reaction kinetics in 1D systems in Subsection 4.1, extends to pattern formation in 2D steady flows in Subsection 4.2, and concludes in Subsection 4.3 by showcasing the method’s scalability in 3D ABC flow dynamics across varying diffusion regimes and reaction models.
All the experiments presented in this work were performed on the HPC2021 system at the University of Hong Kong, which is equipped with 16-core Intel Xeon 6226R processors and an NVIDIA Tesla V100 32GB SXM2 GPU.
@@
The experiments in 1D are conducted on the computational domain $\Omega=[-L,L]$ with $L=60$. The initial condition is given by a uniform distribution on the interval $[0,1]$:
(69) $u_{0}(x)=\chi_{[0,1]}(x)=\begin{cases}1&\text{for }x\in[0,1],\\ 0&\text{otherwise},\end{cases}$
@@
For the reference FDM simulations, we employ a high-resolution discretization with spatial step $\delta x=10^{-3}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions. For the SGIP, we use $N=10^{6}$ particles to approximate the density distribution, with the spatial domain discretized into $K=150$ bins for density estimation, corresponding to a bin width of $\Delta x=0.8$. The time step is set to $\Delta t=0.5$.
We begin by considering three distinct nonlinear reaction terms. The classical FKPP equation, $u_{t}=u_{xx}+u(1-u)$, serves as the baseline. Figure 1 shows that the SGIP results exhibit excellent agreement with the FDM reference, and the wave front propagates with a constant shape, confirming the characteristic traveling wave structure.
Next, a cubic nonlinearity, $u_{t}=u_{xx}+u^{2}(1-u)$, is examined. As shown in Figure 2, this modification leads to a visibly slower propagation speed compared to the standard FKPP case. Finally, we test an Arrhenius-type reaction term, $u_{t}=u_{xx}+e^{-E/u}(1-u)$ with $E=1/2$, which introduces a singularity near $u=0$. Despite this numerical challenge, the SGIP method performs satisfactorily. Figure 3 shows that the wave front propagates at a notably reduced speed while maintaining a coherent profile, further validating the robustness of the method for complex nonlinear reactions.
@@
To quantitatively compare the effects of different reaction terms on propagation speed, we plot the evolution of wave front position over time for all three cases in Figure 4. These numerical experiments comprehensively demonstrate the effectiveness and accuracy of the SGIP method for solving various types of reaction-diffusion equations.
@@
We now extend our investigation to 2D RDA systems to further validate the SGIP method’s capability in handling more complex spatial dynamics. The 2D experiments provide insights into pattern formation, wave propagation in multiple dimensions, and the interaction between diffusion and advection processes.
The 2D simulations are conducted on the computational domain $\Omega=[-60,60]^{2}$, with the initial condition specified by uniformly sampling particle positions from the unit square: $\mathbf{X}_{0}^{i}\sim\text{Uniform}([0,1]^{2})$ for $i=1,\dots,N$, corresponding to an initial density $u_{0}(x,y)=\chi_{[0,1]^{2}}(x,y)$. For the reference FDM simulations, we employ a high-resolution uniform grid with spatial discretization $\delta x=\delta y=10^{-2}$ and temporal step $\delta t=10^{-3}$ to ensure accurate reference solutions. For the SGIP method, $N=3\times 10^{6}$ particles are used to approximate the density distribution, with the spatial domain discretized into $10^{4}$ uniform bins for density estimation and a time step of $\Delta t=0.5$.
@@
Figure 5 demonstrates the evolution of the 2D FKPP equation with uniform initial conditions, showing contour lines at $u=0.1$, $0.5$, and $0.9$. The SGIP method successfully preserves the circular symmetry of the expanding wave front, maintaining rotational invariance in the contour morphology across all three concentration levels.
@@
We examine the interaction between reaction-diffusion dynamics and three representative steady flows through the model $u_{t}=D\Delta u+\mathbf{v}\cdot\nabla u+u(1-u)$. For shear flow $\mathbf{v}=(\sin y,0)$, Figure 6 demonstrates that the SGIP accurately captures the characteristic shear-aligned patterns, showing how the flow stretches and orients the chemical waves. In cellular flow $\mathbf{v}=(-\sin x\cos y,\cos x\sin y)$, Figure 7 reveals how the Lagrangian formulation naturally resolves complex recirculation zones, organizing reaction fronts into circular structures within individual cells, while inducing corrugations along intercellular boundaries due to hyperbolic stagnation points.
For the cat’s eye flow $\mathbf{v}=(-\sin x\cos y,\cos x\sin y)+2(\cos x\sin y,-\sin x\cos y)$, Figure 8 illustrates the competition between diffusion and advection across two regimes: lower diffusivity ($D=0.5$) produces sharp lobe boundaries formed by alternating elliptic islands and hyperbolic channels, while higher diffusivity ($D=1$) enhances cross-channel mixing and smooths front morphology, demonstrating how flow superposition modulates front propagation through competing recirculation and transport regions.
@@
Figure 7: SGIP vs. FDM on cellular flow at $t=20$
@@
Figure 8: SGIP vs. FDM on cat’s eye flow at $t=20$
4.3 3D Numerical Experiments
We now extend our investigation to 3D RDA systems, applying our methods to the ABC flow, which is described by
(70) $u_{t}=D\Delta u+(A\sin z+C\cos y,B\sin x+A\cos z,C\sin y+B\cos x)\cdot\nabla u+u(1-u),$
@@
(a) SGIP with projection at $y=0$ plane
(b) FDM with projection at $y=0$ plane
Figure 9: SGIP vs. FDM on ABC flow with $D=1$ at $t=20$
(a) SGIP with projection at y=0 plane
(b) FDM with projection at y=0 plane
Figure 10: SGIP vs. FDM on ABC flow with $D=0.5$ at $t=20$
@@
The SGIP employs $5\times 10^{6}$ particles with $100^{3}$ spatial bins and a time step $\Delta t=0.5$. Figures 9 and 10 compare both methods on cross-sections through the coordinate planes for diffusion coefficients $D=1$ and $D=0.5$, showing excellent agreement. Since the cross-sectional structures at $x=0$ and $z=0$ are qualitatively similar to those at $y=0$, only the $y=0$ plane is displayed for clarity.
A key advantage of SGIP is its low and stable computational cost. For $D=1$, SGIP requires only 76 s, while the FDM (with $200^{3}$ grids and $\delta t=5\times 10^{-3}$) takes about 337 s. As $D$ decreases to 0.5 (Figure 10), advection and reaction begin to dominate. The SGIP method, using the same parameters ($5\times 10^{6}$ particles, $100^{3}$ bins, $\Delta t=0.5$), maintained a similar computation time, demonstrating its robustness without requiring parameter refinement. In contrast, the FDM required a significantly finer discretization ($300^{3}$ grids, $\delta t=10^{-4}$) to avoid numerical instability, which increased the runtime to approximately $9\times 10^{3}$ s. As $D$ decreases further to 0.1 (Figure 11), the FDM demands an even finer resolution ($400^{3}$ grid points) that becomes computationally prohibitive due to excessive memory requirements on conventional workstations. The SGIP method, however, remains feasible and accurately captures the complex 3D flow structures across all diffusion regimes.
We further validate the SGIP using more complex nonlinear reaction terms (cubic, Arrhenius), maintaining the same computational parameters. The unit ball initial condition is extended to $u_{0}(\mathbf{x})=\chi_{B(0,5)}(\mathbf{x})$ to accommodate the distinct reaction kinetics and ensure clear visualization of the ensuing dynamics.
(a) SGIP with projection at $y=0$ plane
(b) FDM with projection at $y=0$ plane
Figure 12: SGIP vs. FDM on ABC Flow with Reaction Term: $u^{2}(1-u)$ at $t=20$
(a) SGIP with projection at $y=0$ plane
(b) FDM with projection at $y=0$ plane
Figure 13: SGIP vs. FDM on ABC Flow with Reaction Term: $e^{-\frac{1}{2u}}(1-u)$ at $t=20$